Using Gauss’ Theorem calculate the flux of the vector field F = x i^ + y j^ + z k^ through the surface of a cylinder of radius A and height H, which has its axis along the z-axis and the base of the cylinder is on the xy-plane.
1
Expert's answer
2019-04-30T09:44:14-0400

The Gauss theorem states that the flux "\oint_{\partial V} \left( \boldsymbol{F} \cdot \boldsymbol{n} \right) dS" of a vector field "\boldsymbol{F}" over a boundary "\partial V" of volume "V" is equal to the volume integral "\int_V \left( \nabla \cdot \boldsymbol{F} \right) d V". Calculating the divergence of our vector field, we have "\nabla \cdot \boldsymbol{F} = 3", and the volume integral is

"\int_V \left( \nabla \cdot \boldsymbol{F} \right) d V = 3 \int_V d V = 3 V = 3 \pi A^2 H \, ,"


where we have taken into account that "V = \pi A^2 H".


Answer: "3 \pi A^2 H".